Frequency of oscillation of mass m suspended by a spring is n . if length of spring is cut to one third then mass oscillats with n2 . what is relations between n1 and n2??

Grade:12

1 Answers

Himani
6 years ago
T = 2π√(l/g)l1 = l& l2 = l/3so ν = 1/T = (1/2π) x [√g/l) v ∝ 1/√l v2/v1 = √( l1 / l2 ) v2 = √( 3l / l )v1 v2 = (√3)v1